OpenStudy (mary.rojas):

HAPPY BIRTHDAY @missMob <3 Have a wonderful day!!! :D

OpenStudy (missmob):

omg u just posted this in maths

OpenStudy (missmob):

u can get in trouble

OpenStudy (mary.rojas):

ik! just for you! :D yolo

OpenStudy (missmob):

lol thx means alot to me

OpenStudy (mary.rojas):

:D good!

OpenStudy (missmob):

one way to fix this put some maths here

OpenStudy (mary.rojas):

1+1=?

OpenStudy (missmob):

10 + 4 = 14

OpenStudy (missmob):

pie 22/7 or 3.14

OpenStudy (mary.rojas):

lol!

OpenStudy (missmob):

do the fraction of an apple pie

OpenStudy (anonymous):

\[ \begin{array}l\color{red}{\text{H}}\color{orange}{\text{a}}\color{#E6E600}{\text{p}}\color{green}{\text{p}}\color{blue}{\text{y}}\color{purple}{\text{ }}\color{purple}{\text{B}}\color{red}{\text{i}}\color{orange}{\text{r}}\color{#E6E600}{\text{t}}\color{green}{\text{h}}\color{blue}{\text{d}}\color{purple}{\text{a}}\color{purple}{\text{y}}\color{red}{\text{ }}\color{orange}{\text{♫}}\color{#E6E600}{\text{♪}}\color{green}{\text{♬}}\color{blue}{\text{♫}}\color{purple}{\text{♬}}\color{purple}{\text{♩}}\color{red}{\text{}}\end{array} \]

OpenStudy (mary.rojas):

yay! ^

OpenStudy (missmob):



OpenStudy (missmob):

ok

OpenStudy (mary.rojas):

lol!

OpenStudy (missmob):

now fractions like eat one piece 3/4 is the answer eat 2 pieces 2/4 is the answer eat 3 pieces 1/4 is the answer

OpenStudy (missmob):

right ?

OpenStudy (anonymous):